# Unusual critical points between atomic insulating phases

### v1

@@ Abstract
We study a class of quantum phase transitions between featureless bosonic atomic insulators in $(2+1)$ dimensions, where each phase exhibits neither topological order nor protected edge modes. Despite their lack of topology, these insulators may be “obstructed” in the sense that their Wannier centers are not pinned to the physical atomic sites. These insulators represent distinct phases, as no symmetry-preserving adiabatic path connects them. Surprisingly, we find that the critical point between these insulators can host a conformally invariant state described by quantum electrodynamics in $(2+1)$ dimensions (QED3). The emergent electrodynamics at the critical point can be stabilized if the embedding of the microscopic lattice symmetries suppresses the proliferation of monopoles, suggesting that even transitions between trivial phases can harbor rich and unexpected physics. We analyze the mechanism behind this phenomenon, discuss its stability against perturbations, and explore the embedding of lattice symmetries into the continuum through anomaly matching. In all the models we analyze, we confirm that the QED3 is indeed emergeable, in the sense that it is realizable from a local lattice Hamiltonian.
@@ I Introduction
Atomic insulators are typically regarded as the most mundane of quantum phases; such systems are gapped to all excitations, and they lack topological invariants, protected edge modes, or any form of long-range entanglement. Furthermore, their ground states, by definition, can be adiabatically deformed into trivial product states without closing the bulk gap. However, a large body of work has revealed a subtle distinction within this class; while some atomic insulators are trivial in the strictest sense (admitting symmetric, exponentially localized Wannier functions centered at lattice sites), others are “obstructed” Bradlyn et al. (2017); Po et al. (2018); Khalaf et al. (2021); Cano and Bradlyn (2021); Xu et al. (2024). The obstructed atomic insulators are distinguished from the trivial class in that their Wannier centers lie away from the physical lattice sites.
Classes of obstructed atomic insulators (OAIs) have been studied as examples of higher-order topological insulators (HOTIs) Benalcazar et al. (2017a, b); Schindler et al. (2018), where their non-triviality manifests when the system is terminated with boundaries. Unlike symmetry protected topological phases, which host anomalous gapless surface states, OAIs host surface states of lower dimensionality, such as gapless corner modes on a $(2+1)$ dimensional lattice.
@@
We emphasize that the notion of which phase is trivial and which is obstructed is only a relative concept. This is because one must make an arbitrary choice, such as where to place the lattice sites or how to cut the boundary in a finite system, in order to designate a phase as the trivial phase Po et al. (2017a); Bradlyn et al. (2017). One example of this is in the Su–Schrieffer–Heeger (SSH) chain Su et al. (1979), in which the two phases have different polarization values that can be distinguished only relative to each other  [^1] This relative nontriviality is sometimes presented as an absolute one, but any such distinction requires fixing a convention for the unit cell or equivalently, fixing the locations of a background of positive ions. Mathematically, these relative topological distinctions are captured by the mathematical structure of a torsor, not a group or cohomology ring as is usual for the study of topological phases..
This raises a natural question: What can happen at a quantum phase transition between two such “trivial” insulating phases? To this end, we will consider critical points between atomic insulators in $(2+1)d$. A simple example can be realized on the square lattice Benalcazar et al. (2017a, b) with multiple orbitals per site, in which one can tune a phase transition between two insulating states; one phase with localized atomic orbitals that sit on the atomic sites and another phase where the orbitals are centered at each plaquette, as shown in Fig. 1. In this case, at a filling of $n\neq 0\pmod{4}$ particles per unit cell, there is no $C_{4}$ preserving deformation between the two phases. Therefore, there must be a transition between the two insulating phases.
Despite taking place between two trivial phases, the absence of an order parameter highlights that such transitions are not described by Landau–Ginzburg theory. In fact, examples such as deconfined quantum critical points Wang et al. (2017); Senthil (2024) have illustrated that even transitions between conventional ordered phases can be beyond the Landau paradigm, with exotic critical points described by emergent gauge fields and fractionalized excitations Senthil et al. (2004a, b). In this paper, we show that a similar phenomenon occurs between featureless insulators, in which the insulating phases neither break symmetry nor carry topological order.
@@
Going to higher dimension, at the transition between two atomic insulators of bosons in $(2+1)$ dimensions, we find critical points described by $N_{f}=4$ QED3, a theory of $4$ massless Dirac fermions coupled to an emergent $U(1)$ gauge field. This theory has appeared as effective descriptions of systems such as algebraic spin liquids Rantner and Wen (2001); Hermele et al. (2004); Hastings (2000); Hermele et al. (2005, 2007); Ran et al. (2007); Hermele et al. (2008). Furthermore, QED3 can appear as a quantum critical point in strongly correlated lattice magnets Zhang et al. (2025a, b). From analytic and numerical calculations Appelquist et al. (1988); Chester and Pufu (2016); Li (2022a); Albayrak et al. (2022); He et al. (2022); Li (2022b), this theory is believed to be a conformal field theory (CFT) in the IR.
The stability of the QED3 critical point and consequently, the exact nature of the phase transition, is chiefly governed by how the microscopic (UV) lattice symmetries embed into the emergent continuum (IR) symmetries of the quantum critical theory Song et al. (2020, 2019); Zhang et al. (2025a). At the lattice level, space group symmetries such as discrete translations, rotations, and discrete symmetries such as time reversal impose stringent conditions on which operators can be realized in the low energy effective theory. Even independent of the microscopic theory, the embedding of UV symmetries into the IR theory is subject to constraints from the ’t Hooft anomalies of the UV theory, which arise from the Lieb-Schultz-Mattis-Oshikawa-Hastings (LSMOH) theorems Lieb et al. (1961); Oshikawa (2000); Hastings (2004).
@@
The rest of this paper is organized as follows. As a warm-up to our study of atomic insulator transitions, we begin in Sec. II by reviewing the paradigmatic Su-Schrieffer-Heeger (SSH) model Su et al. (1979), both in its fermionic and bosonic forms. This simple one-dimensional model exemplifies how obstructed atomic insulators, though topologically trivial in the absence of symmetry protection, can host distinct phases separated by a critical point and illustrates the key features that will be present in the two-dimensional models. In the fermionic case, the transition involves a Dirac fermion gaining mass, while the bosonic version, analyzed via both exact methods and a parton construction, introduces an emergent gauge field relative to the fermionic case. In both cases, the phase transition can be described by Luttinger liquid theory.
In Sec. III, we generalize the SSH model to two dimensions by analyzing models of atomic insulators on various lattices, such as the Benalcazar-Bernevig-Hughes (BBH) model Benalcazar et al. (2017a, b) on the square lattice. Like the SSH model, each fermionic atomic insulator features Dirac band touchings at the transition. For the bosonic analog, we employ a parton construction that fractionalizes the boson into two fermions coupled to an emergent gauge field, yielding a critical theory described by $N_{f}=4$ QED3. However, the intrinsic bipartiteness of the square and honeycomb lattices introduces a single symmetry-allowed monopole operator that destabilizes the fixed point. Together with a symmetry-allowed mass that tunes the transition, this drives the theory toward confinement and produces a highly multicritical point; we argue the transition is generically first-order. Key to our argument is the observation made in Song et al. (2020) that on a bipartite lattice, $N_{f}=4$ QED3 admits a deformation to $N_{f}=2$ QCD3, from which one can argue there is always at least one monopole operator that is a singlet under all UV symmetries.
In contrast, the $C_{3}$ symmetric breathing kagome lattice is tripartite, and we demonstrate that monopoles are symmetry-disallowed, stabilizing a genuine QED3 critical point. This highlights the strong dependence of criticality in trivial insulator transitions on the underlying (UV) lattice structure.
In Sec. IV, we employ anomaly matching arguments to support our analysis of the critical theories. By matching the LSMOH anomalies for lattice bosons in the IR Zou et al. (2021); Ye et al. (2022), we confirm that all discussed critical theories, even the unstable ones, are realizable by local lattice Hamiltonians. This establishes a rigorous link between the microscopic theory and emergent QED3 critical point. We conclude in Sec. V with broader implications and comments on extensions to three dimensional obstructed atomic insulators.
@@ II.1 Fermionic
A minimal model to illustrate the general case is the SSH model Su et al. (1979) in one dimension, which describes a chain with alternating bonds of two types between atoms, as in polyacetylene. The fermionic SSH chain has a tight binding Hamiltonian
@@
As will also be true for the higher dimensional models, there are many invariants which can be used to sharply distinguish the two phases–these include inversion eigenvalues of the Wannier states at high symmetry $k$ points, the polarization (Zak-Berry phase) Berry (1984); Zak (1989), and even the presence of zero modes when the model is placed on a finite system with open boundary conditions. However, our focus is not on the characterization of each phase; we will instead focus on the critical theory separating the phases.
@@
Once weak, repulsive interactions are added, the critical theory dramatically changes. At the quantum critical point, weak interactions are exactly marginal and lead to a renormalization group (RG) fixed line of (non-Fermi) Tomonaga-Luttinger liquids Haldane (1981); Voit (1995), parametrized by the interaction strength, as shown in Fig. 3. Remarkably, the quasiparticle is lost at the quantum critical point even though both neighboring phases are smoothly connected to band insulators. This is shown in Fig. 4(a-b), where at the critical point, the vanishing of the energy gap is accompanied by a simultaneous vanishing of the quasiparticle residue $Z\rightarrow 0$. Furthermore, the spectral function acquires power-law singularities as $\omega$ approaches the Fermi level, as opposed to being described by a $\delta$-function in the non-interacting limit. In the original fermion Green’s function, this is reflected in the presence of branch cuts instead simple poles. We note the dramatic change of the spectral function from having $\delta$-function behavior to exhibiting a power law singularity in the presence of interactions is special to one-dimensional Luttinger liquids; from standard finite density Fermi liquid theory, in higher dimensions, weak interactions merely broaden the $\delta$-function peak to a Lorentzian and do not destroy the quasiparticle.
@@ II.2 Bosonic
which is simply the fermionic SSH. To characterize the critical point, we can go near the metallic critical point and re-bosonize using standard methods Sénéchal (2004),
@@
The resulting critical theory is $N_{f}=2$ QED2, as there are two Dirac fermions coupled to a $U(1)$ gauge field. The low energy limit of this theory is known Gepner (1985); Affleck (1986); Dempsey et al. (2024) and comprises a massive sector in addition to a gapless sector described by the $SU(2)_{1}$ Wess-Zumino-Witten theory, a CFT with central charge $c=1$. Within abelian bosonization, after bosonizing the fermions $d_{1,2}$ to scalar fields $\phi_{1,2}$, $\phi_{1}+\phi_{2}$ acquires a mass, while $\phi_{1}-\phi_{2}$ remains massless. However, there is no emergent $SU(2)$ symmetry; generically, from microscopic perturbations, there is only a $U(1)\times U(1)$ symmetry, of which a diagonal subgroup is the microscopic charge conservation.
@@ III Phase Transitions between Atomic Insulators in (2+1)d
To study an obstructed atomic insulator transition in two dimensions, we take the BBH model Benalcazar et al. (2017b, a), a representative model of a higher order topological insulator, on the square lattice (Fig. 5). The $C_{4}$ symmetric fermionic model realizes a phase transition between an obstructed atomic insulator with a bulk quadrupole moment and a trivial phase. The obstructed phase exhibits mirror-symmetry protected midgap corner modes on an open geometry.
As mentioned earlier, while this wider class of multipole insulators can be characterized based on the nested Wilson loop formalism Benalcazar et al. (2017b), we will focus only on the critical theories of their phase transitions.
@@
In the bosonic case, fractionalization leads to an emergent QED3 critical point due to the introduction of an emergent $U(1)$ gauge field. However, the proliferation of (renormalization group relevant) symmetry-allowed monopole operators at the critical point leads the critical point to be fine-tuned and ultimately unstable. Importantly, the symmetry-allowed monopole operators arise from the bipartiteness of the original square lattice Song et al. (2020), which enforces a specific embedding of the UV lattice symmetries into the IR that causes the critical point to be fine-tuned. Consequently, the resulting phase transition is likely first order, and we relegate the details of this model and phase transition to Appendix A.
@@
For the analogous atomic insulators on the triangular lattice, we can realize atomic insulators with $C_{6}$ or $C_{3}$ rotation symmetry. The first case is realized by a six-orbital model on the triangular lattice. By viewing the orbitals as their own lattice sites, it can be equivalently viewed as a tight binding model on the breathing honeycomb lattice. The second case is realized by a three-orbital model on the triangular lattice, or equivalently, a tight binding model on the breathing kagome lattice Ezawa (2018). Both lattices are shown in Fig. 6.
@@
In the $C_{3}$ symmetric insulator, however, the lack of bipartiteness allows a possiblity of a stable QED3 critical point and a continuous phase transition. We will explore this model in in the following subsection. We remark that both the $C_{3}$ and $C_{6}$ symmetric models are primitive generators of rotation symmetric higher order topological crystalline insulators in $(2+1)d$ Benalcazar et al. (2019).

### v2

# When Wannier centers jump: Critical points between atomic insulating phases

@@ Abstract
We study a class of quantum phase transitions between featureless bosonic atomic insulators in $(2+1)$ dimensions, where each phase exhibits neither topological order nor protected edge modes. Despite their lack of topology, these insulators may be “obstructed” in the sense that their Wannier centers are not pinned to the physical atomic sites. These insulators represent distinct phases, as no symmetry-preserving adiabatic path connects them. Surprisingly, we find that for certain lattices, the critical point between these insulators can host a conformally invariant state described by quantum electrodynamics in $(2+1)$ dimensions (QED3). The emergent electrodynamics at the critical point can be stabilized if the embedding of the microscopic lattice symmetries suppresses the proliferation of monopoles, suggesting that even transitions between trivial phases can harbor rich and unexpected physics. We analyze the mechanism behind this phenomenon, discuss its stability against perturbations, and explore the embedding of lattice symmetries into the continuum through anomaly matching. In all the models we analyze, we confirm that the QED3 is indeed emergeable, in the sense that it is realizable from a local lattice Hamiltonian.
@@ I Introduction
Atomic insulators are typically regarded as the most mundane of quantum phases; such systems are gapped to all excitations, and they lack topological invariants, protected edge modes, or any form of long-range entanglement. Furthermore, their ground states, by definition, can be adiabatically deformed into trivial product states without closing the bulk gap. However, a large body of work has revealed a subtle distinction within this class; while some atomic insulators are trivial in the strictest sense (admitting symmetric, exponentially localized Wannier functions centered at lattice sites), others are “obstructed” [12, 48, 37, 13, 71]. The obstructed atomic insulators are distinguished from the trivial class in that their Wannier centers lie away from the physical lattice sites.
Classes of obstructed atomic insulators (OAIs) have been studied as examples of higher-order topological insulators (HOTIs) [5, 6, 53], where their non-triviality manifests when the system is terminated with boundaries. Unlike symmetry protected topological phases, which host anomalous gapless surface states, OAIs host surface states of lower dimensionality, such as gapless corner modes on a $(2+1)$ dimensional lattice.
@@
We emphasize that the notion of which phase is trivial and which is obstructed is only a relative concept. This is because one must make an arbitrary choice, such as where to place the lattice sites or how to cut the boundary in a finite system, in order to designate a phase as the trivial phase [46, 12]. One example of this is in the Su–Schrieffer–Heeger (SSH) chain [61], in which the two phases have different polarization values that can be distinguished only relative to each other  [^1] This relative nontriviality is sometimes presented as an absolute one, but any such distinction requires fixing a convention for the unit cell or equivalently, fixing the locations of a background of positive ions. Mathematically, these relative topological distinctions are captured by the mathematical structure of a torsor, not a group or cohomology ring as is usual for the study of topological phases..
This raises a natural question: What can happen at a quantum phase transition between two such “trivial” insulating phases? To this end, we will consider critical points between atomic insulators in $(2+1)d$. A simple example can be realized on the square lattice [5, 6] with multiple orbitals per site, in which one can tune a phase transition between two insulating states; one phase with localized atomic orbitals that sit on the atomic sites and another phase where the orbitals are centered at each plaquette, as shown in Fig. 1. In this case, at a filling of $n\neq 0\pmod{4}$ particles per unit cell, there is no $C_{4}$ preserving deformation between the two phases. Therefore, there must be a transition between the two insulating phases.
Despite taking place between two trivial phases, there is no order parameter description of the phase transition. While two phases with the same microscopic symmetry can, in principle, be separated by a Landau-Ginzburg transition with an emergent scalar order parameter (as in the liquid-gas critical point, where the coarse-grained density is described by a single Ising order parameter), the exotic critical points we explore are not in the universality class of any Landau-Ginzburg theory. Instead, the closing of the single particle gap leads to a metallic critical point described by an interacting gauge theory. In fact, examples such as deconfined quantum critical points [65, 56] have illustrated that even transitions between conventional ordered phases can be beyond the Landau paradigm [57, 55].
In this paper, we show that a similar phenomenon occurs between featureless insulators, in which the insulating phases neither break symmetry nor carry topological order.
@@
Going to higher dimension, at the transition between two atomic insulators of bosons in $(2+1)$ dimensions, we find critical points described by $N_{f}=4$ QED3, a theory of $4$ massless Dirac fermions coupled to an emergent $U(1)$ gauge field. This theory has previously appeared as an effective description of quantum phases of matter such as algebraic spin liquids [52, 31, 27, 32, 33, 51, 30], which can arise as ground states of frustrated magnets. QED3 can also arise as an “unnecessary” quantum critical point within lattice magnets [74, 75]. Moreover, many magnetically ordered states can be obtained by tuning away from a lattice QED3 critical point [60]. From analytic and numerical calculations [3, 16, 38, 2, 29, 39], QED3 is believed to be a conformal field theory (CFT) in the IR. Therefore, in contrast to the theory of deconfined criticality [56], in which later work showed that the putative field theory of the transition was weakly first order and discontinuous [44, 43], the existing evidence for the conformality of $N_{f}=4$ QED3 is very well established.
The stability of the QED3 critical point and consequently, the exact nature of the phase transition, is chiefly governed by how the microscopic (UV) lattice symmetries embed into the emergent continuum (IR) symmetries of the quantum critical theory [59, 60, 74]. At the lattice level, space group symmetries such as discrete translations, rotations, and discrete symmetries such as time reversal impose stringent conditions on which operators can be realized in the low energy effective theory. Even independent of the microscopic theory, the embedding of UV symmetries into the IR theory is subject to constraints from the ’t Hooft anomalies of the UV theory, which arise from the Lieb-Schultz-Mattis-Oshikawa-Hastings (LSMOH) theorems [40, 45, 28].
@@
The rest of this paper is organized as follows. As a warm-up to our study of atomic insulator transitions, we begin in Sec. II by reviewing the paradigmatic Su-Schrieffer-Heeger (SSH) model [61], both in its fermionic and bosonic forms. This simple one-dimensional model exemplifies how obstructed atomic insulators, though topologically trivial in the absence of symmetry protection, can host distinct phases separated by a critical point and illustrates the key features that will be present in the two-dimensional models. In the fermionic case, the transition involves a Dirac fermion gaining mass, while the bosonic version, analyzed via both exact methods and a parton construction, introduces an emergent gauge field relative to the fermionic case. In both cases, the phase transition can be described by Luttinger liquid theory.
In Sec. III, we generalize the SSH model to two dimensions by analyzing models of atomic insulators on various lattices, such as the Benalcazar-Bernevig-Hughes (BBH) model [5, 6] on the square lattice. Like the SSH model, each fermionic atomic insulator features Dirac band touchings at the transition. For the bosonic analog, we employ a parton construction that fractionalizes the boson into two fermions coupled to an emergent gauge field, yielding a critical theory described by $N_{f}=4$ QED3. However, the intrinsic bipartiteness of the square and honeycomb lattices introduces a single symmetry-allowed monopole operator that destabilizes the fixed point. Together with a symmetry-allowed mass that tunes the transition, this drives the theory toward confinement and produces a highly multicritical point; we argue the transition is generically first-order. Key to our argument is the observation made in [59] that on a bipartite lattice, $N_{f}=4$ QED3 admits a deformation to $N_{f}=2$ QCD3, from which one can argue there is always at least one monopole operator that is a singlet under all UV symmetries.
In contrast, the $C_{3}$ symmetric breathing kagome lattice is tripartite, and we demonstrate that monopoles are symmetry-disallowed, making QED3 a natural candidate universality class for a direct continuous transition. This highlights the strong dependence of criticality in trivial insulator transitions on the underlying (UV) lattice structure.
In Sec. IV, we employ anomaly matching arguments to support our analysis of the critical theories. By matching the LSMOH anomalies for lattice bosons in the IR [76, 72], we confirm that all discussed critical theories, even the unstable ones, are realizable by local lattice Hamiltonians. This establishes a rigorous link between the microscopic theory and emergent QED3 critical point. We conclude in Sec. V with broader implications and comments on extensions to three dimensional obstructed atomic insulators.
@@ II.1 Fermionic
A minimal model to illustrate the general case is the SSH model [61] in one dimension, which describes a chain with alternating bonds of two types between atoms, as in polyacetylene. The fermionic SSH chain has a tight binding Hamiltonian
@@
As will also be true for the higher dimensional models, there are many invariants which can be used to sharply distinguish the two phases–these include inversion eigenvalues of the Wannier states at high symmetry $k$ points, the polarization (Zak-Berry phase) [8, 73], and even the presence of zero modes when the model is placed on a finite system with open boundary conditions. However, our focus is not on the characterization of each phase; we will instead focus on the critical theory separating the phases.
@@
Once weak, repulsive interactions are added, the critical theory dramatically changes. At the quantum critical point, weak interactions are exactly marginal and lead to a renormalization group (RG) fixed line of (non-Fermi) Tomonaga-Luttinger liquids [26, 64], parametrized by the interaction strength, as shown in Fig. 3. Remarkably, the quasiparticle is lost at the quantum critical point even though both neighboring phases are smoothly connected to band insulators. This is shown in Fig. 4(a-b), where at the critical point, the vanishing of the energy gap is accompanied by a simultaneous vanishing of the quasiparticle residue $Z\rightarrow 0$. Furthermore, the spectral function acquires power-law singularities as $\omega$ approaches the Fermi level, as opposed to being described by a $\delta$-function in the non-interacting limit. In the original fermion Green’s function, this is reflected in the presence of branch cuts instead simple poles. We note the dramatic change of the spectral function from having $\delta$-function behavior to exhibiting a power law singularity in the presence of interactions is special to one-dimensional Luttinger liquids; from standard finite density Fermi liquid theory, in higher dimensions, weak interactions merely broaden the $\delta$-function peak to a Lorentzian and do not destroy the quasiparticle.
@@ II.2 Bosonic
which is simply the fermionic SSH. To characterize the critical point, we can go near the metallic critical point and re-bosonize using standard methods [54],
@@
The resulting critical theory is $N_{f}=2$ QED2, as there are two Dirac fermions coupled to a $U(1)$ gauge field. The low energy limit of this theory is known [25, 1, 20] and comprises a massive sector in addition to a gapless sector described by the $SU(2)_{1}$ Wess-Zumino-Witten theory, a CFT with central charge $c=1$. Within abelian bosonization, after bosonizing the fermions $d_{1,2}$ to scalar fields $\phi_{1,2}$, $\phi_{1}+\phi_{2}$ acquires a mass, while $\phi_{1}-\phi_{2}$ remains massless. However, there is no emergent $SU(2)$ symmetry; generically, from microscopic perturbations, there is only a $U(1)\times U(1)$ symmetry, of which a diagonal subgroup is the microscopic charge conservation.
@@ III Phase Transitions between Atomic Insulators in (2+1)d
To study an obstructed atomic insulator transition in two dimensions, we take the BBH model [6, 5], a representative model of a higher order topological insulator, on the square lattice (Fig. 5). The $C_{4}$ symmetric fermionic model realizes a phase transition between an obstructed atomic insulator with a bulk quadrupole moment and a trivial phase. The obstructed phase exhibits mirror-symmetry protected midgap corner modes on an open geometry.
As mentioned earlier, while this wider class of multipole insulators can be characterized based on the nested Wilson loop formalism [6], we will focus only on the critical theories of their phase transitions.
@@
In the bosonic case, fractionalization leads to an emergent QED3 critical point due to the introduction of an emergent $U(1)$ gauge field. However, the proliferation of (renormalization group relevant) symmetry-allowed monopole operators at the critical point leads the critical point to be fine-tuned and ultimately unstable. Importantly, the symmetry-allowed monopole operators arise from the bipartiteness of the original square lattice [59], which enforces a specific embedding of the UV lattice symmetries into the IR that causes the critical point to be fine-tuned. Consequently, the resulting phase transition is likely first order, and we relegate the details of this model and phase transition to Appendix A.
@@
For the analogous atomic insulators on the triangular lattice, we can realize atomic insulators with $C_{6}$ or $C_{3}$ rotation symmetry. The first case is realized by a six-orbital model on the triangular lattice. By viewing the orbitals as their own lattice sites, it can be equivalently viewed as a tight binding model on the breathing honeycomb lattice. The second case is realized by a three-orbital model on the triangular lattice, or equivalently, a tight binding model on the breathing kagome lattice [23]. Both lattices are shown in Fig. 6.
@@
In the $C_{3}$ symmetric insulator, however, the lack of bipartiteness allows a possiblity of a stable QED3 critical point and a continuous phase transition. We will explore this model in in the following subsection. We remark that both the $C_{3}$ and $C_{6}$ symmetric models are primitive generators of rotation symmetric higher order topological crystalline insulators in $(2+1)d$ [7].